A student eats a dinner containing 8.0x10 raised to power 6j of energy. He wishes to do an equivalent amount of work in a nearby gym by lifting a 60kg object.

How many times must he raise the object to expand this much energy? Assume that he raises it a distance of 2.0m each time.


1
Expert's answer
2021-04-06T13:50:29-0400
"W=mghn,""n=\dfrac{W}{mgh}=\dfrac{8\cdot10^6\ J}{60\ kg\cdot9.8\ \dfrac{m}{s^2}\cdot2\ m}=6800\ times."